a car accelerates from rest at a constant rate of 2m/s2 for some time.Then it retards at a constant rate of 4m/s2 and comes to rest.If it remains in motion for 3 seconds,the total distance covered by the car is??
  • 2m
  • 3m
  • 4m
  • 6m
PLS EXPLAIN IT FAST

Grade:12th Pass

1 Answers

Ayushi Verma
6 years ago
because car starts from rest thats why initial velocity vi=0& given acceleration ai= 2m/s^2let fibal velocity vf = v in time tinow, vf =vi + aiti so, v = 0 + 2ti so, v = 2ti so, ti = v/2---------(i)then it again coming to rest and its initial velocity is that which it attained when it started from rest is v thats why when it again coming to rest at that period vi=v, vf=0,aii = -4m/s^2 & time be tii so,we can write as so, vf = vi + aii*tiiso,0 = v - 4tiiso tii = v/4 ---------(ii)adding equations (i) & (ii) ti + tii = v/2 +v/4so, 3= 3v/4so, v =4m/s^2